Determine the finite population correction factor for each of the following. (a) N-2000 and n 1500 (c) N = 2000 and n = 150 (e) What happens to the finite population correction factor as the sample size n decreases but the population size N remains the same? (b) N- 2000 and n-200 (d) N= 2000 and n = 100 (a) The finite population correction factor is (Round to three decimal places as needed.) (b) The finite population correction factor is Round to three decimal places as needed.) (c) The finite population correction factor is (Round to three decimal places as needed.) (d) The finite population correction factor is (Round to three decimal places as needed.) (e) What happens to the finite population correction factor as the sample size n decreases but the population size N remains the same? A. The finite population correction factor approaches 1. B. The finite population correction factor approaches 0. ° C. The finite population correction factor does not approach any set value. ? D. The finite population correction factor approaches 0.707 2
Answer #1

as finite population correction factor =sqrt((N-n)/(N-1))

a)

finite population correction factor =sqrt(500/1999)=0.500

b)finite population correction factor =sqrt(1800/1999)=0.949

c))finite population correction factor =sqrt(1850/1999)=0.962

d))finite population correction factor =sqrt(1900/1999)=0.975

e)

option A